OpenStudy (anonymous):

find the roots of 2(x-3)^2-11=0 to two decimal places by isolating (x-3)^2 then taking the square root of both sides

OpenStudy (anonymous):

To isolate the (x-3)^2, add 11 to both sides and then divide by 2. \[(x-3)^{2}=11/2\] Now take the square root of both sides. FInally add 3 to both sides. Remember, when you take the square root, you will have plus/minus in front of the radical.

OpenStudy (anonymous):

x= (11/2)^1/2 +3 now use calculator

OpenStudy (anonymous):

x=5.35

OpenStudy (anonymous):

OMG! you rock!

OpenStudy (anonymous):

The problem statement asks for "the roots". What is the value of the other root?

OpenStudy (anonymous):

it should be\[\pm\sqrt{11/2}+3\] You need both the positive and negative radical. Plug both into calculator.

OpenStudy (anonymous):

Correct,\[-\sqrt{11/2}+3 = 0.654792 \]

OpenStudy (anonymous):

THANKS! :D